OpenStudy (anonymous):

a3b3c -------- a-3b-3c-1

OpenStudy (pfenn1):

Is this the correct expression?\[\frac{a^3b^3c}{a^{-3}b^{-3}c^{-1}}\]

OpenStudy (anonymous):

yes

OpenStudy (jackellyn):

a exponent 3-(-3)=6 b exponent 3-(-3) =6 c exponent 1-(-1)=2 So, \[a ^{6}b ^{6}c ^{2}\]

OpenStudy (pfenn1):

Recall that \[\large \frac{x^a}{x^b}=x^{a-b}\]